# Designing Adaptive User Interfaces for mHealth applications targeting chronic disease: A User-Centric Approach

### v1

@@ 4.5.1. Trade-off between user burden, user support and controllability
( , , ) Users have consistently highlighted the importance of having various options for controlling and supporting adaptations. However, effectively managing the complexity of providing users with the appropriate level of control and support remains a challenge that can affect the usability and effectiveness of AUIs (Paymans et al., 2004; Bunt et al., 2004; Eiband et al., 2021). User control over adaptation can yield significant benefits by empowering users to tailor their experiences to better suit their needs and preferences (Parra and Brusilovsky, 2015; Kay, 2001). Additionally, user-controlled adaptation fosters a sense of ownership and agency, leading to increased engagement and motivation to utilise the system effectively (Sundar and Marathe, 2010; Sundar, 2008). Our study participants expressed a preference for control mechanisms such as dedicated adaptation dashboards, selective system component adaptation, and configurable adaptation factors. However, our findings also highlight the importance of considering users’ competence levels, the potential cognitive load, and the risk of adaptation processes becoming intrusive (Kurtz, 2005). Allowing excessive control may lead to distraction and inefficiency, particularly among users who lack the necessary knowledge or interest to make informed decisions (Jameson and Schwarzkopf, 2002) ( ). Furthermore, as discussed by Kay (2001), user preferences for control can vary significantly. Some users may feel obligated to have control over the adaptation and other elements of the user interface, while others may be less inclined towards control (Bunt et al., 2004). Providing users with an excessive number of control options can result in distraction and fatigue when making decisions ( ). Additionally, differences in user perspectives on controllability may also be related to specific tasks user performs, as highlighted by Gajos and Chauncey (2017). This relationship, focusing on task-related characteristics, is further explored in the study by Lavie and Meyer (2010). They found that the proportion of routine and non-routine tasks, as well as the efforts required for certain tasks, can affect the actual performance of AUIs. This underscores the importance of considering these factors, as users often engage in a mix of tasks when using mHealth apps, with some tasks being more frequent or demanding than others (Bunt et al., 2004). In tasks that are essential and frequently performed, the importance of user controllability diminishes, and users may prefer the system to handle these tasks without requiring explicit confirmation or manual adjustments (Peissner and Edlin-White, 2013). Conversely, for tasks that require more effort, users might expect to have greater control over the adaptation. Therefore, delegating control to users requires considering various factors, including the nature of the task and user characteristics, rather than focusing solely on one specific aspect.
Apart from the controllability, research has attempted to support the adaptation process. For example, some studies have explored the use of animated transitions to demonstrate the adaptation process to users (Dessart et al., 2011), while others have investigated the effectiveness of providing detailed explanations (Kühme, 1993). However, it seems questionable whether these methods should be applied to adaptations that attempt to compensate for users’ perceptual, cognitive, and motor impairments. The provision of support materials may not necessarily reduce cognitive demands, as additional support itself can impose an additional burden on users (Kulesza et al., 2013). Meanwhile, users may not always find value in support materials, such as explanations, especially when they lack control over the adaptation process, perceive that the system is functioning effectively, or consider the time and effort required to understand the information outweighing the benefits (Bunt et al., 2012). Therefore, while enhancements in controllability and support are beneficial, it is crucial to carefully consider their implementation and ensure that they meet users’ actual needs ( ).
@@ 4.5.2. Trade-off for usability issues
( , , ) AUIs are increasingly seen as a solution to cope with the growing diversity of usage contexts, devices, and users (Browne et al., 1990; Höök, 1998; Norcio and Stanley, 1989). They offer solutions to various usability issues in mobile applications, including improving accuracy, efficiency, and user learning, as well as addressing information overload and helping in the use of complex systems (Browne et al., 1990; Höök, 1998). However, previous research on AUIs has also demonstrated a trade-off between adaptive mechanisms and usability (Paymans et al., 2004; Jameson, 2007; Mezhoudi et al., 2015; Deuschel, 2018; Höök, 2000; Glass et al., 2008; Bunt et al., 2004; Gajos et al., 2006, 2008; Turk et al., 2019). We found several usability challenges associated with AUIs, including privacy concerns, predictability issues, comprehensibility difficulties, and UI obtrusiveness in our user study. Interestingly, our study reveals that certain strategies proposed by participants to enhance the adaptation design can alleviate or exacerbate some existing usability challenges. For example, aligning visual elements and icons with user preferences can reduce UI obtrusiveness ( ). However, it is crucial to note that introducing icons like lifelike characters, such as a doctor icon or an ambulance icon, may create unrealistically high expectations regarding the system’s competence in understanding and adapting to the user (Jameson, 2007; Sproull et al., 1996). Creating the right expectations in users is crucial for effective design, ensuring neither overestimation nor underestimation of the system’s competence (Höök, 2000). Mismatches in expectations can directly reduce trust and subsequently lead to a decreased intention to use the interface again (Glass et al., 2008; Pu and Chen, 2006). Other strategies, such as attempts to enhance control or privacy over the adaptation, may necessitate additional user interactions for proposed adaptations, data inputs to extract feedback for the generated adaptation, or system notifications for changes made through adaptations. This could potentially increase the risk of disruption and divert user attention from their previous focus (Jameson, 2007; Deuschel and Scully, 2016) ( ).
We have observed that while certain strategies can mitigate specific usability challenges, they may inadvertently exacerbate others. Meanwhile, the effectiveness of a strategy depends on its appropriate application, as it can have varying impacts on usability. Despite the initial advantages of a strategy aimed at addressing a specific usability challenge, its excessive application can lead to the negation of potential benefits and the emergence of new usability issues (Jameson, 2007). For example, aligning with chronic disease guidelines and providing explanatory materials can improve adaptation comprehensibility ( , ). However, an excessive provision of support content can become obtrusive, difficult to grasp, and impair user learnability (Eiband et al., 2021; Paymans et al., 2004). Similarly, step-by-step adaptation at regular intervals could improve UI predictability and could also increase the intrusiveness of the adaptation process ( ). However, constantly adapting the UI can be obtrusive and disrupt the workflow of users (Wesson et al., 2010). Future research efforts should aim to improve the benefits offered by AUIs while mitigating potential challenges associated with them.
@@
Mitigating specific usability challenges in AUI design often entails trade-offs with other usability goals, emphasising the intricate nature of achieving a balance between adaptation and usability, and necessitating efforts to minimise conflicting demands. The key lies in striking a balance between the various demands of an interface. Our findings highlight two key considerations to deal with conflicting demands.
@@
1) User priorities. Users often have varying priorities regarding conflicting usability goals (Rich and McGee, 2004). For example, in our user study, we observed that some participants preferred interfaces with consistent layouts and design elements across various screens, as it fosters familiarity and ease of use. In contrast, others prioritise adaptability and enjoy exploring new interface designs and features, even if it means sacrificing some consistency. Similarly, users may have different privacy priorities. While some individuals prioritise privacy and are cautious about sharing personal information, others are more inclined to offer data in return for customised user interfaces. These examples highlight the critical importance of offering alternative solutions that accommodate the diverse usability priorities of users (Bunt et al., 2004; Gajos et al., 2008). This approach assumes that different users may be open to trade-offs among different usability aspects. Harper et al. (2015) found that users reported usability issues with a system but remained overall satisfied because the benefits of having control options outweighed the encountered inconveniences.
@@
2) Granularity. Usability objectives such as predictability, comprehensibility and controllability can be achieved at different levels of granularity (Bunt et al., 2004; Jameson, 2007). Our observations revealed that making minor adjustments during the adaptation process could not significantly impact the overall user experience. Instead, small incremental changes could alter the user’s familiarity with the UI, leading to a less consistent and less predictable user experience (Gajos et al., 2006; Bunt et al., 2004). Additionally, these low-level granularity adaptations may have a limited impact on the comprehensibility of the adaptations, as they typically affect only small and limited parts of the UI (Abrahão et al., 2021). High-level granularity adaptation involves modifying multiple aspects of the UI simultaneously (Cristea and Calvi, 2003; Zeidler et al., 2013), such as altering the overall colour scheme, font styles, navigation menus, and placement of key features at the same time. It may introduce more usability challenges, as it could be perceived as a completely new app from the user’s perspective. Typically, high-level granularity adaptations occur less frequently and may occur only once to tailor the app to the user’s specific needs, which could compensate for some usability issues. This approach is particularly advantageous when users first engage with the system, as they do not have preconceived expectations of its appearance or functionality (Bunt et al., 2004). In such scenarios, users approach the system without investing time or effort into learning its interface or forming expectations about its design and functionality. Therefore, any adaptation that occurs may not significantly affect the predictability or comprehensibility of the UI, as users lack a pre-established reference point for comparison. However, it may pose challenges for users who are less familiar with such applications and lack insights into which design and functionality would best suit their needs (Gajos et al., 2006).
Finding the right balance between these factors is essential for designing AUIs that effectively manage the trade-off between adaptation and usability. Some usability issues introduced by AUIs can be mitigated without necessarily improving the user’s mental model of the adaptive system (Paymans et al., 2004; Rosman et al., 2014). However, systems that actively help users overcome usability issues tend to foster a deeper understanding and trust in the system’s ability to adapt effectively to their needs, encouraging sustained engagement and usage of mobile health apps (Martin-Hammond et al., 2019). As users become more familiar with the adaptive capabilities and benefits of the system, they are likely to develop a deeper appreciation of its role in enhancing their overall experience and well-being.
@@ 4.5.3. Trade-off between independence and assistance
( , ) User support typically focuses on facilitating the execution of app functions, but this type of support may hinder user learning process and diminish their overall experience with the system (Lanier, 1995). Learning involves acquiring skills, knowledge, and competencies within a specific domain, which allows individuals to evolve and become more independent. However, this learning process often entails significant costs, including the time and effort spent acquiring and executing activities, which can be error-prone and time-consuming ( ). It can pose a significant challenge for users when they are required to learn to use the system each time it undergoes adaptation. In addition to providing user support, mitigating the cognitive effort required to adapt the system can be achieved by allowing other users, such as caregivers or health professionals, to manage the adaptation process. For example, a chronic disease management app could be tailored by caregivers or health professionals, albeit at the expense of the user’s independent utilisation of the technology ( ). The decision to invest time in domain-specific learning or delegate tasks to the system should ultimately be made by the user. Offering users a spectrum of options between independent and assisted app usage may still be beneficial. The design of support features should align with users’ objectives and contextual factors (Fischer, 2006), accommodating users’ preferences and comfort levels in embracing challenges across various situations.
@@
The inevitability of trade-offs complicates decision-making in UI adaptation, as users frequently differ in their prioritisation of conflicting goals, level of competence, objectives for using the app, and the specific usage scenario. It is imperative to provide alternative solutions that cater to users with varying priorities (Billsus and Pazzani, 2007). Furthermore, exploring innovative UI design solutions can help mitigate the need for users to choose suboptimal options. For example, Cheverst et al. (2005) developed an office control system prototype that initially offered users two task delegation options. Despite these choices, users were dissatisfied with the usability of the prototype. However, acceptance markedly improved when designers expanded the design space by introducing a separate small screen for the office control system.
5. Findings from survey
5.1. Survey participants
Concurrently with our focus groups and interviews, we surveyed 90 participants with chronic diseases, all of whom have experience using mHealth applications. Their detailed demographic information is summarised in Table 4. Most of the participants identified as men (56%), between the ages of 18 and 74. In particular, the age group of 25-34 constituted the largest group, comprising about 33% of the total participants. There was representation from older age groups, with 16% of participants over 55 years of age and 21% between 35 and 44 years of age. Geographically, our survey captured responses from a diverse set of countries, with approximately half of the responses originating from Australia. China and the USA are also significant contributors, accounting for 23% and 13% of the responses, respectively. In terms of educational background, a bachelor’s degree is the most common attainment (44%), followed by 34% with education levels below a bachelor’s degree and 21% with higher degrees, such as master’s or doctorate degrees.
We grouped participant chronic health conditions into four broader groups: cardiometabolic (e.g., diabetes, high blood pressure, obesity, and heart disease), respiratory (e.g. allergies, asthma, and chronic lung disease), immune-related (e.g., cancer, Parkinson’s disease, and compromised immune system), and mental health conditions (Camacho-Rivera et al., 2020). Cardiometabolic diseases were the most prevalent, reported by 52% of the participants. It is worth noting that some participants reported multiple chronic diseases, such as having both diabetes and high blood pressure, or a combination of cardiometabolic disease and another type of disease, such as asthma (a respiratory disease). Indicates a phenomenon known as multimorbidity (Islam et al., 2014), where individuals may experience two or more chronic conditions simultaneously. This is also consistent with other studies (Anderson et al., 2016). Participants were asked to indicate their familiarity with mHealth apps. As shown in Table 4, respondents reported using various types of mHealth apps, with health-promoting and self-monitoring apps being the most prevalent, utilised by 84% of participants. These applications typically cater to functions related to fitness, medication, and diet, which is consistent with findings from previous research (Kayyali et al., 2017). This prevalence can be attributed to the ongoing need for individuals with chronic diseases to monitor their health status and adhere to prescribed treatment regimens (Hamine et al., 2015). The primary motivations cited for using mHealth apps include symptom monitoring (58%), promoting physical activity (54%), and managing dietary intake (47%). Participants were also asked about the frequency of using the mHealth app (see Figure 8). In our survey study, 54% of participants reported using health apps daily, while 33% reported weekly usage. Most users spend between 1 to 20 minutes per session, aligning with findings from prior research (Robbins et al., 2017; Kc et al., 2021). Users who access the app two or more times per day typically allocate a shorter duration per session. On the contrary, people who use the app less frequently (less than once a month or less than once a week) tend to engage in longer sessions. The former group may use the app for quick, frequent interactions or brief check-ins throughout the day, while the latter may delve into more comprehensive interactions during less frequent usage, possibly for specific tasks or content consumption.
| Demographics | # | % of Participants |
| Age of the Participants1 | ||
| 18-24, 17/19% — 25-34, 30/33% — 35-44, 19/21% — 45-54, 10/11% — 55-64, 11/12% — 65-74, 3/3% | ||
| Gender of the Participants2 | ||
| Female, 38/42%, — Male, 50/56%, — Prefer not to say, 2/3% | ||
| Countries of the Participants | ||
| Australia | 44 | 49% |
| China | 21 | 23% |
| USA | 12 | 13% |
| UK | 6 | 7% |
| Other (Nigeria, Canada, Korea, Spain and Sri Lanka) | 7 | 8% |
| Education information of the Participants3 | ||
| Less than Bachelor’s degree | 31 | 34% |
| Bachelor’s degree | 40 | 44% |
| Postgraduate (Master’s and Doctoral degree) | 19 | 21% |
| Categories of chronic disease4 | ||
| Cardiometabolic | 47 | 52% |
| Immune-related | 31 | 34% |
| Mental health conditions | 7 | 8% |
| Respiratory | 11 | 12% |
| Type of mHealth application used 5 | ||
| Health promoting and self-monitoring application | 76 | 84% |
| Informative application | 41 | 46% |
| Assistive application | 29 | 32% |
| Communication application | 24 | 27% |
| Health game application | 16 | 18% |
| Rehabilitation application | 16 | 18% |
| Exercise application | 7 | 8% |
| Main reason to use the app | ||
| To monitor my chronic disease symptoms | 52 | 58% |
| Increase my physicial activity levels | 49 | 54% |
| To track what I eat | 42 | 47% |
| To get education about my chronic disease | 34 | 38% |
| Help with weight loss | 28 | 32% |
| Manage my medications | 26 | 29% |
@@ 5.1. Survey participants
| 1,2,3: The percentage does not strictly add up to 100% due to rounding. 4: The percentage does not strictly add up to 100% due to multimorbidity. 5: The percentage does not strictly add up to 100% due to the use of multiple applications. | ||
@@
| Different types of adaptation | Data source of adaptation | ||||
@@
| Different data collection method | Preferred level of involvement in the adaptation | ||||

### v2

@@ 4.2.1. Trade-off between user burden, user support and controllability
( , , ) Users have consistently highlighted the importance of having various options to control and support adaptations. However, effectively managing the complexity of providing users with control and support remains a challenge (Paymans et al., 2004; Bunt et al., 2004; Eiband et al., 2021). User control over adaptation offers significant benefits, empowering users to customize their experiences to better meet their needs and preferences (Parra and Brusilovsky, 2015; Kay, 2001). It also encourages a sense of ownership and agency, boosting engagement and motivation for effective system use (Sundar and Marathe, 2010; Sundar, 2008). Leaving excessive control can lead to distraction and inefficiency, particularly among users who lack the necessary knowledge or interest to make informed decisions (Jameson and Schwarzkopf, 2002) ( ). Furthermore, as discussed in Kay (2001), the user preferences for control can vary significantly. Variations in user preferences for controlling adaptations and interface elements can be explained by Hofstede’s cultural dimensions model (Hofstede, 2001), which suggests that users of individualistic cultures often prefer personal control, while those from collectivist cultures can rely on preset options (Bunt et al., 2004). Too many control options can lead to distraction and fatigue ( ). Furthermore, differences in user perspectives on controllability may also be related to specific tasks the user performs, as highlighted by Gajos and Chauncey (2017). This underscores the importance of considering these factors, as users often engage in a mix of tasks when using mHealth applications, and some tasks are more frequent or demanding than others (Bunt et al., 2004). In essential and frequent tasks, user controllability becomes less important, and users may prefer the system to manage these tasks automatically (Peissner and Edlin-White, 2013). Therefore, delegating control to users requires considering various factors, including cultural influence and the nature of the task. In addition to controllability, research has attempted to support the adaptation process. For example, some studies have explored the use of animated transitions to demonstrate the adaptation process to users (Dessart et al., 2011), while others have investigated the effectiveness of providing detailed explanations (Kühme, 1993). However, the provision of support materials may not necessarily reduce cognitive demands, as additional support itself can impose an additional burden on users (Kulesza et al., 2013). Users may not always value support materials such as explanations, especially when they lack control over the adaptation process, perceive the system as effective, or find the effort to understand the information not worth the benefits (Bunt et al., 2012). Therefore, while improvements in controllability and support are beneficial, it is crucial to carefully consider their implementation and ensure that they meet user needs ( ).
@@ 4.2.2. Trade-off for usability issues
( , , ) AUIs are increasingly seen as a solution to cope with the growing diversity of usage contexts, devices, and users (Browne et al., 1990; Höök, 1998; Norcio and Stanley, 1989). They offer solutions to various usability issues in mobile applications, including improving accuracy, efficiency, and user learning, as well as addressing information overload and helping in the use of complex systems (Browne et al., 1990; Höök, 1998). However, previous research on AUIs has also shown a trade-off between adaptive mechanisms and usability (Paymans et al., 2004; Jameson, 2007; Mezhoudi et al., 2015; Deuschel, 2018; Höök, 2000; Glass et al., 2008; Bunt et al., 2004; Gajos et al., 2006, 2008; Turk et al., 2019). We found several usability challenges associated with AUI, including privacy concerns, predictability issues, comprehensibility difficulties, and UI obtrusiveness in our user study. Interestingly, our study reveals that certain recommendations proposed by participants to improve the adaptation design can alleviate or exacerbate some existing usability challenges. For example, aligning visual elements and icons with user preferences can reduce the obtrusiveness of the UI ( ). However, introducing lifelike character icons, such as a doctor or ambulance, can unrealistically increase expectations of the system’s competence in adapting to the user (Jameson, 2007; Sproull et al., 1996). Setting accurate user expectations is key for effective design, as misaligned expectations can lower trust and reduce the likelihood of reusing the interface, emphasizing the importance of avoiding both overestimation and underestimation of the system competence (Höök, 2000; Glass et al., 2008; Pu and Chen, 2006). Recommendations such as improving adaptation control or privacy can require additional user interactions, feedback input, or system notifications, potentially disrupting and distracting users (Jameson, 2007; Deuschel and Scully, 2016) ( ). We have observed that while certain recommendations can mitigate specific usability challenges, they can inadvertently exacerbate others (Jameson, 2007). Meanwhile, the effectiveness of a recommendation depends on its appropriate application, as it can have varying impacts on usability. For example, aligning with chronic disease and providing explanatory materials can improve adaptation comprehensibility ( , ). However, excessive provision of support content can become obtrusive, difficult to grasp, and alter user learnability (Eiband et al., 2021; Paymans et al., 2004). Similarly, step-by-step adaptation at regular intervals could improve UI predictability and could also increase the intrusiveness of the adaptation process ( ). However, constantly adapting the UI can be obtrusive and disrupt user workflow (Wesson et al., 2010).
@@
Mitigating specific usability challenges in AUI design often involves trade-offs with other usability goals. Our findings highlight two key considerations for dealing with conflicting demands.
@@
1) User priorities. Users often have varying priorities regarding conflicting usability goals (Rich and McGee, 2004). For example, in our user study, we observed that some participants preferred interfaces with consistent layouts and design elements across various screens, as it fosters familiarity and ease of use. In contrast, others prioritize adaptations and enjoy exploring new interface designs and features, even if it means sacrificing some consistency. Similarly, users may have different privacy priorities. Although some people prioritize privacy and are cautious about sharing personal information, others are more inclined to offer data in return for a customized UI. These examples highlight the critical importance of offering alternative solutions that accommodate the diverse usability priorities of users (Bunt et al., 2004; Gajos et al., 2008). This approach assumes that different users may be open to trade-offs among different usability aspects. Harper et al. (2015) found that users reported usability problems with a system but were overall satisfied because the benefits of having control options outweighed the inconveniences encountered.
@@
2) Granularity. Usability objectives such as predictability, comprehensibility, and controllability can be achieved at different levels of granularity (Bunt et al., 2004; Jameson, 2007). Our observations revealed that making minor adjustments during the adaptation process could not significantly impact the overall user experience. Furthermore, low-level granularity adaptations have a minimal effect on comprehensibility, as they only affect small parts of the UI (Abrahão et al., 2021). The adaptation of high-level granularity involves modifying multiple aspects of the UI simultaneously (Cristea and Calvi, 2003; Zeidler et al., 2013), such as altering the overall color scheme, font styles, navigation menus, and placement of key features at the same time. It may introduce more usability challenges, as it could be perceived as a completely new application from the user’s perspective. High-level granularity adaptations are infrequent and may happen just once to customize the application for the user, helping to address some usability issues. This approach is particularly advantageous when users first engage with the system, as they do not have preconceived expectations of its appearance or functionality (Bunt et al., 2004). However, it may pose challenges for users who are less familiar with such applications and lack insight into which design and functionality would best suit their needs (Gajos et al., 2006). Some usability issues introduced by AUI can be mitigated without necessarily improving the user’s mental model of the adaptive system (Paymans et al., 2004; Rosman et al., 2014). Systems that help users overcome these problems foster trust and understanding, encouraging the continued use of mHealth applications (Martin-Hammond et al., 2019). As users recognize the system’s adaptive benefits, they appreciate its role in enhancing their experience and well-being.
@@ 4.2.3. Trade-off between independence and assistance
( , ) User support typically focuses on facilitating the execution of application functions, but this type of support may hinder the user learning process and diminish their overall experience with the system (Lanier, 1995). Learning is the process of acquiring skills, knowledge, and competencies in a specific domain, allowing greater independence. However, this comes with costs such as time and effort, and can be prone to errors and time-consuming ( ). Mitigating the cognitive effort to adapt the system can be achieved by having caregivers or health professionals manage the process ( ). Users should decide whether to invest time in learning or delegate tasks to the system or caregivers. Providing options for independent and assisted application use can be advantageous. Support features should align with user goals and context, accommodating their preferences and comfort with challenges (Fischer, 2006).
@@
The inevitability of trade-offs complicates decision making in UI adaptation, as users frequently differ in their prioritization of conflicting goals, level of competence, objectives for using the app, and the specific usage scenario. It is imperative to provide alternative solutions that cater to users with varying priorities (Billsus and Pazzani, 2007).
5. Stage One: Findings of the user survey
5.1. User Survey Participants
| Demographics # % of Participants Age1 18-24 17 19% 25-34 30 33% 35-44 19 21% 45-54 10 11% 55-64 11 12% 65-74 3 3% Gender2 Female 38 42% Male 50 56% Prefer not to say 2 3% Education3 Less than Bachelor’s degree 31 34% Bachelor’s degree 40 44% Postgraduate 19 21% Chronic Disease Categories4 Cardiometabolic 47 52% Immune-related 31 34% Mental health conditions 7 8% Respiratory 11 12% | Demographics # % of Participants Country of Residence Australia 44 49% China 21 23% USA 12 13% UK 6 7% Other (Nigeria, Canada, Korea, Spain, Sri Lanka) 7 8% Main reason to use the app5 To monitor my chronic disease symptoms 52 58% Increase my physicial activity levels 49 54% To track what I eat 42 47% To get education about my chronic disease 34 38% Help with weight loss 28 32% Manage my medications 26 29% Type of mHealth Application Used5 Health promoting and self-monitoring 76 84% Informative application 41 46% Assistive application 29 32% Communication application 24 27% Health game 16 18% Rehabilitation application 16 18% Exercise application 7 8% |
@@ 5.1. User Survey Participants
| * 1,2,3: The percentage does not strictly add up to 100% due to rounding. 4: The percentage does not strictly add up to 100% due to multimorbidity. 5: The percentage does not strictly add up to 100% due to the use of multiple applications/purpose. | |
Concurrently with focus group and interview studies, we surveyed 90 participants with chronic diseases, all of whom have experience using mHealth applications. Their detailed demographic information is summarized in Table 1. Most of the participants identified as men (56%), between the ages of 18 and 74. In particular, the 25-34 age group constituted the largest group, comprising about 33% of all participants. There was representation from older age groups, with 16% of participants over 55 years of age and 21% between 35 and 44 years of age. Geographically, our survey captured responses from a diverse set of countries, with approximately half of the responses coming from Australia. China and the USA are also significant contributors, accounting for 23% and 13% of the responses, respectively. In terms of educational background, a bachelor’s degree is the most common attainment (44%), followed by 34% with education levels below a bachelor’s degree and 21% with higher degrees, such as master’s or doctorate degrees.
The chronic disease reported by the participants were systematically categorized into four groups: cardiometabolic (e.g., diabetes, high blood pressure, obesity, and heart disease), respiratory (e.g., allergies, asthma, and chronic lung disease), immune-related (e.g., cancer, Parkinson’s disease, and compromised immune system), and mental health conditions (Camacho-Rivera et al., 2020). Cardiometabolic diseases were the most prevalent, reported by 52% of the participants. It should be noted that some participants reported multiple chronic diseases, such as having diabetes and high blood pressure, or a combination of cardiometabolic disease and another type of disease, such as asthma (a respiratory disease). Indicates a phenomenon known as multimorbidity (Islam et al., 2014), where individuals may experience two or more chronic diseases simultaneously, a finding that is also consistent with other studies (Anderson et al., 2016). Participants were asked to indicate their familiarity with the mHealth applications. As shown in Table 1, respondents reported using various types of mHealth applications, with health-promoting and self-monitoring applications being the most prevalent, utilized by 84% of participants. These applications typically target functions related to fitness, medication, and diet, which is consistent with findings from previous research (Kayyali et al., 2017). This prevalence can be attributed to the ongoing need for individuals with chronic diseases to monitor their health status and follow prescribed treatment regimens (Hamine et al., 2015). The primary motivations cited for using mHealth applications include symptom monitoring (58%), promoting physical activity (54%), and management of dietary intake (47%). Participants were also asked about the frequency of using the mHealth app. In our survey study, 54% of the participants reported using health applications daily, while 33% reported weekly usage. Most users spend between 1 to 20 minutes per session, aligning with findings from previous research (Robbins et al., 2017; Kc et al., 2021). Users who access the application two or more times per day typically allocate a shorter duration per session. In contrast, people who use it less frequently (less than once a month or less than once a week) tend to engage in longer sessions. The former group may use the application for quick, frequent interactions or brief check-ins throughout the day, while the latter may delve into more comprehensive interactions during less frequent usage, possibly for specific tasks or content consumption.
@@
| Different types of adaptation* | Data source of adaptation* | ||||
@@
| Different data collection method* | Preferred level of involvement in the adaptation | ||||
@@
| *: Because respondents could select multiple options, the percentages do not sum precisely to 100%. | |||||
5.2. Different Types of Adaptations
In the earlier SLR on AUIs for chronic disease-related applications (Wang et al., 2023b) compiled a comprehensive taxonomy of adaptation implemented by researchers to support users in managing chronic diseases. Our aim with this survey is to discern the users’ perspectives regarding the significance and value attributed to these adaptations. Our survey revealed a diversity of preferences among the respondents, with no singular adaptation type dominating over others. Content complexity (59%) and graphic design (58%) emerged as the most prevalent, closely followed by the rearrangement of the interface elements (47%) and multimodal interaction (46%). In contrast, the sound effects exhibited the lowest frequency among the listed adaptations (28%) (see Table 2). This finding is not consistent with the dominant trends in existing AUI studies in the domain of chronic diseases (Wang et al., 2023b), where the SLR emphasized graphic design (presentation adaptation), as the predominant adaptation type; however, the survey findings highlighted the significant emphasis of users on content adaptation and some types of behavior adaptation. It highlights the significance of synchronizing researchers’ initiatives with user needs and priorities to guarantee the efficient design and application of AUIs in chronic disease management tools.